An automobile of mass 103kg moves with a velocity of 60m/s. The driver applies the brakes to reduce the velocity to 20m/s. What is the energy dissipated during the application of the brakes?
1
Expert's answer
2021-08-11T09:54:22-0400

Solution.

"m=103kg;"

"v_1=60m\/s;"

"v_2=20m\/s;"

"\Delta W=\dfrac{mv_1^2}{2}-\dfrac{mv_2^2}{2};"

"\Delta W=\dfrac{103\sdot60^2}{2}-\dfrac{103\sdot20^2}{2}=164800J;"

Answer: "\Delta W=164800J."